### v1

# Strong asymptotic giant branch stars’ spectral features in distant quiescent galaxies: Impact on galaxy evolution

Abstract
Age-dating and weighting stellar populations in galaxies at various cosmic epochs are essential steps to study galaxy formation through cosmic times. Evolutionary population synthesis models with different input physics are used towards this aim. In particular, the contribution from the thermally pulsing asymptotic-giant-branch (TP-AGB) stellar phase, which peaks for intermediate-age 0.6–2 Gyr systems, has been debated upon for decades. Here we report the detection of strong cool star signatures in the rest-frame near-infrared spectra of three young (1 Gyr), massive () quiescent galaxies at large look-back time, z=1–2, using JWST/NIRSpec. The co-existence of oxygen- and carbon-type absorption features, spectral edges and features from rare species such as Vanadium, and possibly Zirconium, reveal a strong contribution from TP-AGB stars. Population synthesis models with significant TP-AGB contribution reproduce the observations considerably better than those with weak TP-AGB, which are those commonly used. These findings call for revisions of published stellar population fitting results, pointing to lower masses and younger ages, with additional implications on cosmic dust production and chemical enrichment. These results will stimulate new generations of improved models informed by these and future observations.
School of Astronomy and Space Science, Nanjing University, Nanjing, 210093,China
Université Paris-Saclay, Université Paris Cité, CEA, CNRS, AIM, Paris, 91191,France
Key Laboratory of Modern Astronomy and Astrophysics (Nanjing University), Ministry of Education, Nanjing, 210093, China
Institute of Cosmology and Gravitation, University of Portsmouth, Dennis Sciama Building, Burnaby Road, Portsmouth PO1 3FX, UK
NSF’s National Optical-Infrared Astronomy Research Laboratory, 950 N. Cherry Ave. Tucson, AZ 85719, USA
Instituto de Física, Pontificia Universidad Católica de Valparaíso, Casilla 4059, Valparaíso, Chile
INAF-Osservatorio Astronomico di Padova, Vicolo dell’Osservatorio 5, I-35122 Padova, Italy
University of Massachusetts Amherst, 710 North Pleasant Street, Amherst, MA 01003-9305, USA
Department of Astronomy, The University of Texas at Austin, Austin, TX, USA
INAF-Osservatorio Astronomico di Roma, via di Frascati 33, 00078 Monte Porzio Catone, Italy
Department of Physics and Astronomy, University of California, 900 University Ave, Riverside, CA 92521, USA
Instituto de Astrofísica de Canarias (IAC), 38205 La Laguna, Tenerife, Spain
Space Telescope Science Institute, 3700 San Martin Dr., Baltimore, MD 21218, USA
Universidad de la Laguna, La Laguna, Tenerife, Spain
Université Paris-Cité, LERMA - Observatoire de Paris, PSL, Paris, France
Laboratory for Multiwavelength Astrophysics, School of Physics and Astronomy, Rochester Institute of Technology, 84 Lomb Memorial Drive, Rochester, NY 14623, USA
Department of Physics, University of California Merced, 5200 Lake Road, Merced, CA 95343, USA
Department of Physics and Astronomy, Texas A&M University, College Station, TX, 77843-4242, USA
George P. and Cynthia Woods Mitchell Institute for Fundamental Physics and Astronomy, Texas A&M University, College Station, TX, 77843-4242, USA
ESA/AURA Space Telescope Science Institute
Department of Physics and Astronomy, University of California, Los Angeles, CA, 90095, USA
Kavli Institute for Cosmology, University of Cambridge, Madingley Road, Cambridge, CB3 0HA, UK
Cavendish Laboratory, University of Cambridge, 19 JJ Thomson Avenue, Cambridge, CB3 0HE, UK
1 Introduction
The TP-AGB – an advanced evolutionary stage of intermediate-mass stars (1-10) – is a potentially essential component of stellar population models. The study of the TP-AGB contribution to the light of stellar population has a long history. The first estimate goes back over 30 years [1] and was based on the fuel consumption theorem, thus predicting that the TP-AGB contribution to the bolometric light of a simple stellar population (SSP) could approach or even exceed at intermediate ages from Myr to a few Gyr. Subsequent models estimated a contribution to up to 80% of the integrated near-infrared (NIR) emission of stellar populations with ages in the range of 0.2 2 Gyr [2, 3, 4, 5, 6].
However, the actual energetic contribution and spectral composition of this phase are highly controversial, with different models adopting distinctively different prescriptions for both the energetics and the stellar spectra during this phase [7, 3, 8, 9]. These uncertainties stem from two complications typical of the TP-AGB phase, namely its double-shell burning regime leading to instabilities and ‘thermal pulses’ occurring on short timescales alongside a strong mass-loss, both requiring calibration of free parameters such as convection, mixing and mass loss with observational data [10, 11]. As a consequence, stellar evolutionary computations, which are the backbone of evolutionary population synthesis models, typically terminate at the end of the Early Asymptotic Giant Branch (E-AGB) phase, and the inclusion of the TP-AGB in integrated models is performed separately either using specific TP-AGB tracks [8, 4] or fuel-consumption-theorem estimated TP-AGB energetics [3].
Historically, the early Maraston models (i.e. M05 [3]) based on the fuel consumption theorem [12] pioneered the incorporation of the TP-AGB phase into population synthesis models by calibrating the energetics of TP-AGB tracks with observed photometric data and adopting observed oxygen-rich (O-rich) and carbon-rich (C-rich) spectra [13, 14]. These M05 models, forecasting the onset of the TP-AGB phase as a sharp transition (e.g., a rapid increase in the TP-AGB bolometric and near-IR contribution to the energetics of stellar populations, see Figure 3 in Ref. [2]) around 0.2 Gyr and lasting up to 2 Gyr, are referred to as “TP-AGB-heavy”. A similar modelling approach is found in the Conroy models (i.e., C09 [9]), but with two modifications, namely a substantial reduction of the energetics and a shift towards older ages (1-3 Gyr) for the relevant population epoch, which dilute the TP-AGB ‘phase transition’. Therefore, the C09 models are referred to as “TP-AGB-light”. Following the availability of new data for the calibration and adopting some of the C09 arguments, the calibration of the TP-AGB contribution as a function of age was revised using Magellanic Clouds globular clusters data [10], based on which an update of the M05 models was presented (so-called M13). The M13 models have a reduced TP-AGB fuel consumption and a slightly older onset age (600 Myr) with respect to M05, but they still forecast a sharp transition for the TP-AGB onset and a sizable fuel consumption. The M13 models are referred to as “TP-AGB-mild”. On a different route, the widely used Bruzual & Charlot models (i.e., BC03 [7]), based on isochrone synthesis of stellar evolutionary tracks [15], adopt different energetics and spectra for the TP-AGB phase. Since this results in a negligible TP-AGB contribution to the integrated spectrum, the BC03 models are referred to as “TP-AGB-poor”.
A number of studies have attempted to observationally constrain the contribution of the TP-AGB phase by fitting spectra and photometric spectral energy distributions (SEDs) of different kinds of galaxies (post-starbursts, Seyferts, spirals, and passive) over a range of redshifts, but with conflicting results [16, 17, 18, 19, 20, 21]. It was early suggested (M05) that high-redshift galaxies with low star formation (‘quiescent’) are the ideal laboratories for this test, as their mass-weighted ages (1 Gyr) correspond to the expected maximum possible contribution of the TP-AGB stars to the rest-frame NIR spectrum of such galaxies. High quality observations of NIR galaxy spectra dominated by TP-AGB stars are still missing: ground-based spectra below 2.3m are impacted by atmospheric emission and absorption, and HST NIR grism spectra lack sufficient sensitivity. The JWST has now the wavelength coverage and sensitivity to finally settle this issue.
2 Results
2.1 Selection of quiescent candidates
We selected 1084 quiescent galaxy candidates at using a catalog [22] in the EGS CANDELS field [23, 24], based on the UVJ diagram [25] and SED modeling (adopting exponentially declining star formation history, namely SFH and requiring ). The limit imposed on the sample ensures that the age-sensitive spectral region around rest-frame 4000Å break is covered by JWST/NIRSpec with good sensitivity. We cross-matched these phometrically-selected candidates with the 1491 galaxies with secure spectroscopic redshift observed with JWST/NIRSpec during the Cosmic Evolution Early Release Science (CEERS) program [26, 27] and during a subsequent NIRSpec follow-up program (DD-2750) [28]. This resulted in a spectroscopic sample of 31 galaxies, for which we visually inspected the spectra. Some spectra have low signal-to-noise ratios (SNR) because the galaxies are too faint or were poorly centered in the NIRSpec’s micro-shutter apertures. Other spectra exhibit strong emission lines consistent with either active star formation or an active galactic nucleus (AGN), or that have redshifts that are too high () to provide adequate coverage of the relevant rest-frame NIR wavelengths. We retained three galaxies, IDs = 8595, 9025, and D36123 with redshifts at , whose spectra clearly show the D4000 break, no obvious emission lines, and also have relatively high signal-to-noise ratio (SNR). This is a small yet representative sample fulfilling these criteria.
2.2 The spectrum of D36123
Galaxy D36123, the primary target in this work, exhibits an exceptionally high-quality PRISM spectrum with an average of SNR 187 per spectral pixel (Fig.1; plotted as a double histogram to show the noise). Its relatively bright flux (mag=21.4 at 3.6m), steep radial surface brightness profile (Sérsic 4, see Table 3), longer exposure time and accurate centering into the micro-shutters, which minimizes the effect of path loss and aperture correction (see Methods), are the reasons for the remarkable quality of this spectrum.
A large number of features are clearly detected at rest-frame wavelengths from 0.3 to 2.0 m, for which we propose identifications (see labels in Fig.1). Besides common absorption features at Å typical of quiescent galaxies, such as Ca H+K, H, H+G4300 and absorption from heavy elements (such as Mg, Ca, and Fe), which have been observed in quiescent galaxies’ spectra at high redshift [29], the rest-frame NIR spectrum exhibits numerous absorption features and band edges. The relatively low spectral resolution (R30–300) of the NIRSpec PRISM helps in the detection of these broad features. The proposed identifications are obtained by comparing our spectrum with empirical libraries of C-rich [14] and O-rich [13] TP-AGB stars (see Fig.2). These observed spectral features are related to diverse stellar types in the AGB phase, i.e. E-AGB and TP-AGB [30, 31, 32, 33]. The deep CN edges in Fig.2a (most striking is the CN1.1m edge, see Fig.1, where the continuum suddenly drops down by 15% in the observed spectrum) are particularly enhanced in rich N-type carbon stars [30, 31, 32], that are produced by the third dredge-up and therefore demonstrate the presence of a contribution by TP-AGB stars. On the other hand, the TiO and VO absorption bands (see Fig.2b) are produced in the spectra of M-type stars in either the RGB or AGB phases [31]. However, the best-fit spectrum of BC03 models, that includes only the RGB contribution, is almost featureless in the NIR (Fig. 3) and we conclude that also the oxygen-rich features in the spectrum of D36123 are due to TP-AGB stars. This is the first time that multiple TiO absorptions concomitant with and features are identified in an observed galaxy spectrum, pointing to the simultaneous presence of O-rich and C-rich cool stars. This is naturally expected in presence of a broad distribution of stellar metallicities, which must exist in massive galaxies, such that the metal-poor stars evolve preferentially into carbon-rich TP-AGB stars, and the metal-rich ones into oxygen-rich (M Type) TP-AGB stars. Since V and Zr are only produced by s-process nucleosynthesis, those VO and ZrO features (see Fig.2b) are expected to be detectable only in the coolest TP-AGB stars [31, 33], demonstrating the prominence of the TP-AGB phase in the light emitted by this quiescent galaxy. Note also the broad depressions due to water, again indicating the presence of O-rich M-type stars.
We then performed spectral fitting on the observed spectrum of D36123, employing four models – BC03, C09, M13, and M05 (spanning the full range from TP-AGB poor to heavy, see Fig.3 and Table 1). We used delayed exponentially declining SFH with a large variety of timescales and stellar ages, interpolated over a grid of metallicities, and allowed for dust extinction. In convolving the models to match the resolution of observed spectra, we considered the instrumental resolution, the wavelength sampling of the spectral template models, and the estimated stellar velocity dispersion of the galaxy (see Methods). We find that our model fitting results are robust against (1) more complex star-formation histories allowing for bursts and sudden truncations; (2) linear combinations of models with different ages and metallicities (which is equivalent to non-parametric SFH); (3) alternative attenuation curves or without considering the reddening by dust; (4) variations to the aperture correction recipe; (5) modifications of the spectral extraction apertures over 3 or 5 pixels; (6) modelling of emission lines. The minimum reduced from each model is barely affected when these variations are considered.
The “TP-AGB-mild/heavy” (i.e., M13 and M05) models can produce the largest fraction of the detected features, while the “TP-AGB-poor/light” predict much weaker spectral features in the NIR, as Fig.3 illustrates. Among the four models, the TP-AGB-mild model, M13, provides by far the best fit. Additionally, M13 also identifies the majority of TP-AGB features (i.e., four TiO absorptions and most CN edges). However, there are strong spectral features that are clearly not present (or much weaker than in the observed spectrum) even in the M13 best-fit spectrum, like several TiO and C absorptions. The strong CN1.1m spectral edge is only present in the M13/M05 models. Interestingly, the MgII-band at m, part of Lick set of indexes [34] is much weaker in all models than in the data, pointing to a strong contribution from TiO bands all the way to 0.5m, as expected for M-type stars (and, possibly, to [/Fe]-enhancement [35]). The M05 shows the worst overall fit due to clear discrepancy at (see Fig.10 in Methods), but still displays significantly more numerous matching TP-AGB features in the NIR than the C09 and BC03 models (see Fig.3), noticeably the long-sought C absorption around 1.75m. Moreover, when fitting the spectrum over the restricted rest-frame wavelength range Å all models converge to a mass-weighted age of about 1 Gyr, which is indeed the age where TP-AGB features are expected to be at peak [3, 10]. Therefore, we can conclude that it is the treatment of the TP-AGB in the models that most affects the goodness of fit in our quiescent galaxies. Even for the best-fit model M13, the minimum reduced is large, indicating that also this model is lacking some stellar ingredients that are present in the galaxy, and are needed.
Some of the observed features are also found in the coolest phase of (M-type) red-giant-branch (RGB) stars and in the red-supergiant (RSG) phase of massive stars, with ages around 10 Myr [13, 30, 33]. We could safely exclude that RSG stars contribute substantially due to the inferred mass-weighted age (1 Gyr) and negligible residual star formation rate, when fitting only the optical spectrum (see Table 1). This is suggestive of a negligible contribution from short-lived massive stars. Instead, M-type RGB stars certainly contribute to the NIR spectrum at these ages, to a degree that is, however, model-dependant because the onset and development of the RGB phase depends on the overshooting assumed in stellar tracks [36, 3]. The full RGB contribution to the total light in the K-band around 1 Gyr is according to M05 models and for BC03-type models and the upper part of the RGB where M-giants form is a further fraction of this. Therefore, while contributing, M-type RGB stars do not dominate the spectrum. We verified this point by comparing our observed spectrum to a further stellar population model including empirical spectra for the RGB (and other phases; E-MILES [37]), but not including a treatment of the TP-AGB phase. This model cannot reproduce more features than the BC03/C09 models.
The best-fitting physical parameters presented in Table 1 exemplify the impact of the chosen model on the derived properties. For instance, comparing results between BC03 and M13 in the full rest-frame spectral range, there is only 30% discrepancy in stellar mass, but the impact on ages is strong with 1 Gyr (the M13 age is almost three times lower than the BC03 age – all reported age measurements in this work are mass-weighted), despite the higher metallicity preferred by the BC03 fit. These age differences affect the derived formation epochs and evolutionary timescales of galaxies. The BC03 best-fit age of 1.7 Gyr is consistent with a RGB contribution of about for those models (see Figure 2 in Ref.[38]), which likely contributes to some of the observed features and NIR flux. However, this fitted age is still too young for the contribution from red giant stars to be dominant, further supporting the arguments described above.
2.3 The spectra of galaxies 8595 and 9025
The spectra of the other two fainter quiescent galaxies, 8595 (mag=24 at 3.6m, observed with the PRISM) and 9025 (mag=23 at 3.6m, observed with the medium-resolution gratings but rebinned to the PRISM resolution), have substantially lower SNR of 15 and 7. Nevertheless, the SNR is still sufficient for the direct detection of individual absorption features in object 8595 (Fig.4), some of the strongest also observed in object 9025 (Fig.5). This includes some TiO absorptions and CN spectral edges, and tentatively a C absorption.
Table 2 presents the results of the best spectral fits of objects 8595 and 9025, based on the different models. For object 8595, we carefully account for a neighbouring bright galaxy when measuring its photometry to estimate spectral aperture corrections, following a similar approach to what is done for object D36123 (see Methods). Again, M13 yields the best-fit spectral fits for both galaxies, with 0.8 for object 8595, and 1.7 for object 9025. The derived stellar masses differ at most by dex between the M13 and BC03 models, but the best-fit ages are even more strongly dependent on the specific model used (stellar age differences from 0.1 to 1 Gyr). In Table 2, the values denote the difference with respect to the M13 model. The small cumulative probability suggests the preference for the M13 model is statistically significant at the ( confidence level) in the overall spectral fit for these two galaxies.
3 Discussion
Overall, we conclude that the spectra of distant quiescent galaxies in this study show clear evidence for a substantial contribution of TP-AGB stars of both C and M type, at variance with what predicted by TP-AGB-poor and -light models. The models with significant TP-AGB contribution, such as M13 or M05, are in better agreement with the NIR features in the spectra of these high-redshift quiescent galaxies, and TP-AGB-mild models (i.e., M13) can fit their entire spectra better. These support our conclusion that the TP-AGB is the primary model ingredient driving the fitting of our spectrum.
Previous work[18] based on high-quality NIR spectra of local spiral/Seyfert galaxies reached similar conclusions, favouring M05 models over BC03 models because of the TP-AGB component. Those local spectra cover a wide wavelength range similar to our high-redshift data and display a variety of features, several of which can only be attributed to TP-AGB stars. Previously, a detection of the CN 1.1m band in the NIR spectra of local AGNs was interpreted as a signature of intermediate age populations evolving along the TP-AGB phase [32]. Although these local observations are remarkable, JWST spectroscopy can directly observe massive, high-redshift galaxies, with low spread in stellar generations, for their stars to evolve to the TP-AGB phase in sync at ages 1 Gyr when the TP-AGB energetic contribution is maximum [3, 5]. This maximizes the contribution of the short TP-AGB phase (3 Myr for an individual star) to the integrated galaxy SED, with respect to systems with ongoing star formation and a lower mass fraction of stars in the appropriate age range. The dominance of the TP-AGB phase also implies that the NIR spectrum will display all its complexity and thus the SED fit becomes challenging.
Indeed, we find that even for the best performing M13 model the is large for the high SNR spectrum of D36123 that shows stronger absorption features than the model. By separating optical (m) and NIR (m), we confirmed that it is the fit in the NIR spectrum leading to high values of , since we find a minimum reduced –11 in the optical spectrum for the four different models (C09 performing best in this case), while all reduced values remain large (40–60) when fitting only the NIR spectrum (see Table 1). More in detail, it appears that the rest-frame m range, where the strongest TiO features are present, is the most problematic region (see Table 1 and Fig.11 in the Methods) since when we exclude this range, both optical and NIR can be fitted to better than 2% accuracy by the M13 model (Fig.11; but not as well by BC03 and CO09).
Several model ingradients could be responsible for this discrepancy. First of all, M13 and M05 use empirical spectra for just the TP-AGB phase. For the other phase that strongly contribute to the NIR, the RGB, classical Kurucz-type theoretical spectra from model atmospheres are adopted, which - as it is well known - do not adequately match all NIR absorption features [39, 40]. Similarly, models based on the empirical RGB spectra perform better, but not for all features [39]. Notably, studies by different researchers show that all CO NIR spectral features from H to K in local massive early-type galaxies, whose NIR spectrum is dominated by RGB stars, are systematically stronger than those in the models based on empirical libraries [41]. This is also supported by the analysis of observed NIR spectra from local spirals and star-forming galaxies [42]. All this suggests that D36123 contain also stars that are more metal rich than those included in the model.
Our model comparison suggests that extensions to the best-fitting models are needed to improve the match to massive galaxy spectra. The M13 models adopt local samples of C-rich and O-rich spectra, which may not cover the range of metallicity and abundance ratios hosted in distant massive ellipticals. The metallicity dependence in M13 models is included using a theoretical prescription [12], by which more metal-rich models have more fuel consumption in O-rich rather than in C-rich phases, but still using the same observed spectra. In particular, M13 models do not treat the relative contributions of C-rich and O-rich TP-AGB stars as a free parameter. Moreover, the calibration of stellar population models requires samples of stellar generations that are as simple as possible, i.e., star clusters, with available independent age and chemical composition from Color Magnitude Diagram (CMD) fitting and stellar spectroscopy. For the TP-AGB phase these calibrating clusters also need to span a wide age range, extending down to Myr. Star clusters with these properties exist in the Magellanic Clouds and have traditionally served to calibrate stellar population models [2, 43, 10]. The limitation is that these clusters have sub-solar metallicity and possibly solar or sub-solar element ratios.
Another issue is a galaxy specific chemical pattern. It is well-known that massive galaxies, both locally [35, 44] as well as at high redshift [29, 45, 46], feature [/Fe]-enhanced stellar populations. The same galaxies are also found to host enhanced ratios of [C/Fe] and [N/Fe] [47, 48] and possibly also in Na [49, 46, 50]. Empirical libraries based on local, solar-scaled RGB spectra, and the population models based on them, may not reflect these chemical patterns, causing discrepancies. This would mirror a long-standing discrepancy between optical spectral features of massive galaxies and early population models [51], which was solved by -enhanced population models [52]. Galaxy spectra like those we present here, with young ages and super-solar metallicity, and possibly various non-solar element enhancements, are unique, with no known counterpart in the local universe. They challenge the models, but also offer the invaluable opportunity to calibrate them, with far-reaching benefit also for stellar evolution and model atmosphere theories.
Astronomers depend on evolutionary population synthesis models to infer the timeline of galaxy formation and evolution. In turn, the model age scale depends on the input stellar evolution in terms of stellar tracks (and their input) and the energetics of stellar phases. We derive systematically younger ages and lower masses when fitting these spectra with TP-AGB-mild models (i.e., M13) versus TP-AGB-poor/-light models (i.e., BC03/C09), especially at high redshift. The age differences range from 100 Myr to 1 Gyr (see Tables 1 and 2). Younger galaxy ages can have implications for the inferred physics of galaxy evolution, such as gas accretion and star formation quenching. Also galaxy simulations adopt evolutionary population synthesis models as input, hence our results have an impact on the predicted spectra and photometry of simulated galaxies found in cosmological and other simulations [6]. In that case, the age is fixed within the simulation, but the simulated galaxy SEDs and the calibrations through comparisons to observations are obviously affected. The AGB stellar phase has also a major impact on galaxies dust production and chemical evolution particularly for heavy elements [53]. A higher fuel for this phase, as supported by our findings, has implications on guiding the relevant calculations for these processes [54], thus potentially having an even broader impact on galaxy evolution.
The spectra presented here demonstrate the presence of TP-AGB spectral features, including those from carbon stars, in the integrated NIR light from high-redshift, quiescent galaxies. These features are first seen in their richness in galaxy spectra. Clearly, our very small galaxy sample does not allow us to infer the statistical impact of the TP-AGB emission on the derived stellar population properties for young quiescent galaxies in general. Previous photometric studies [5, 19] have shown that, on average, fitting distant galaxies using models with heavier TP-AGB contributions yields younger ages and lower stellar masses than TP-AGB-poor models, but also with significant galaxy-by-galaxy differences. The usual degeneracies between age, reddening and metallicity still apply to some extent, and for individual galaxies, the SED fitting results can depend on unknown details of the past star formation histories (for example). From our small sample, it is also difficult to evaluate how frequently these TP-AGB features are present in early galaxy spectra. NIRSpec observations for a larger, homogeneously selected sample of galaxies can address these questions and help us to quantify the broader consequences of TP-AGB emission for determining galaxy physical parameters. Such observations may also help guide further improvement of stellar population models.
| Rest-frame | Property | BC03 | C09 | M13 | M05 |
| spectral range | TP-AGB-poor | TP-AGB-light | TP-AGB-mild | TP-AGB-heavy | |
| RA = 14h : 19m : 34.258s | DEC = +52 : 56 : 23.079 | ||||
| redshift | 1.074 | 1.075 | 1.082 | 1.075 | |
| ] | 1.493 | 1.489 | 1.167 | 1.552 | |
| age [Gyr] | 1.661 | 0.751 | 0.621 | 2.224 | |
| full | 1.944 | 2.000 | 1.589 | 0.982 | |
| (0.3–2m) | Av | 0.000 | 0.156 | 0.424 | 0.000 |
| [Gyr] | 0.241 | 0.104 | 0.100 | 0.707 | |
| SFR | 0.294 | 0.099 | 0.311 | 1.320 | |
| SFR/SFR | 0.019 | 0.014 | 0.024 | 0.368 | |
| 59.6 | 52.9 | 39.0 | 102.6 | ||
| redshift | 1.078 | 1.083 | 1.078 | 1.078 | |
| ] | 1.330 | 1.406 | 1.140 | 1.151 | |
| age [Gyr] | 0.961 | 0.810 | 0.909 | 0.969 | |
| optical | 1.525 | 2.000 | 2.183 | 1.002 | |
| (m) | Av | 0.068 | 0.000 | 0.000 | 0.000 |
| [Gyr] | 0.007 | 0.091 | 0.001 | 0.007 | |
| SFR | 0.000 | 0.093 | 0.000 | 0.000 | |
| SFR/SFR | 0 | 0 | 0 | 0 | |
| 11.4 | 9.0 | 11.2 | 11.2 | ||
| redshift | 1.074 | 1.080 | 1.083 | 1.080 | |
| ] | 1.679 | 1.514 | 1.247 | 0.556 | |
| age [Gyr] | 1.961 | 0.783 | 0.490 | 0.207 | |
| NIR | 1.883 | 2.000 | 0.692 | 0.500 | |
| (m) | Av | 0.000 | 0.103 | 0.706 | 0.982 |
| [Gyr] | 0.497 | 0.057 | 0.023 | 2.605 | |
| SFR | 0.984 | 0.000 | 0.000 | 23.725 | |
| SFR/SFR | 0.207 | 0 | 0 | 0.200 | |
| 60.7 | 58.5 | 40.7 | 66.6 | ||
| redshift | 1.076 | 1.093 | 1.081 | 1.075 | |
| ] | 1.581 | 1.750 | 1.076 | 1.535 | |
| age [Gyr] | 1.627 | 1.446 | 0.712 | 2.183 | |
| full | 2.300 | 0.300 | 1.285 | 0.990 | |
| (excluding | Av | 0.000 | 0.078 | 0.068 | 0.000 |
| 0.5–1m) | [Gyr] | 0.028 | 0.127 | 0.095 | 0.690 |
| SFR | 0.000 | 0.000 | 0.084 | 1.306 | |
| SFR/SFR | 0 | 0 | 0.011 | 0.363 | |
| 32.2 | 23.1 | 16.3 | 83.7 | ||
The 1st column denotes the rest-frame range of wavelength adopted in the spectral fit. The 3rd to 6th columns list the best-fitting physical properties derived by four models with different TP-AGB treatments. Specifically, the total stellar masses are based on the Chabrier[55] IMF. The best-fitting parameters and errors are derived from the distribution of , with the former obtained at -minimisation, and the errors determined after rescaling to . SFR/SFR is a measure of quiescence, where the SFR is derived from the best fit by adopting delay- models, and the SFR is corresponding to the SFR maximum at .
| Target | Property | BC03 | C09 | M13 | M05 |
| (Rest-frame) | TP-AGB-poor | TP-AGB-light | TP-AGB-mild | TP-AGB-heavy | |
| RA = 14h : 20m : 42.240s | DEC = +53 : 01 : 50.298 | ||||
| redshift | 1.739 | 1.746 | 1.746 | 1.741 | |
| ] | 0.791 | 0.736 | 0.545 | 0.676 | |
| age [Gyr] | 1.913 | 0.798 | 0.732 | 1.908 | |
| 8595 | 1.711 | 2.000 | 1.584 | 1.878 | |
| (full) | Av | 0.012 | 0.357 | 0.368 | 0.013 |
| [Gyr] | 0.283 | 0.134 | 0.133 | 0.403 | |
| SFR | 0.073 | 0.049 | 0.061 | 0.232 | |
| SFR/SFR | 0.021 | 0.042 | 0.061 | 0.113 | |
| 0.821 | 0.950 | 0.794 | 0.935 | ||
| ) | 0.00103 | 0 | – | 0 | |
| RA = 14h : 19m : 33.991s | DEC = +52 : 51 : 56.290 | ||||
| redshift | 1.545 | 1.546 | 1.569 | 1.569 | |
| ] | 0.759 | 0.859 | 0.755 | 0.746 | |
| age [Gyr] | 1.560 | 1.580 | 0.272 | 0.182 | |
| 9025 | 2.123 | 0.300 | 1.700 | 1.140 | |
| (full+HST) | Av | 0.351 | 0.397 | 1.125 | 1.337 |
| [Gyr] | 0.304 | 0.316 | 0.019 | 0.008 | |
| SFR | 0.181 | 0.230 | 0.048 | 0.027 | |
| SFR/SFR | 0.082 | 0.092 | 0 | 0 | |
| 1.702 | 1.838 | 1.634 | 1.649 | ||
| ) | 0.00001 | 0 | – | 0.03767 | |
For both objects, the spectral fit utilizes the entire rest-frame wavelength range. Additionally, to better constrain the age-sensitive wavelength around 4000Å, two additional HST bands (F606W+F814W) are employed simultaneously in the fit of object 9025 (see Methods). Measurements and errors are as in Tab.1, except that we did not rescale the reduced to 1. is computed relative to the M13 model.
4 Methods
4.1 NIRSpec observation and reduction
In this study, we present NIR spectra for three quiescent galaxies, IDs=D36123, 8595, and 9025. We cross-matched the catalog in CANDELS EGS field [22] with the NIRSpec detections in the DD-ERS-1345 (CEERS, PI: Steven Finkelstein) and DD-2075 (PI: Pablo Arrabal Haro) programs. Both D36123 and 8595 were observed with the PRISM [56], spanning 0.60-5.30m with varying spectral resolution R 30 at 1.2m to R at 5m. Object 9025 was observed with the G140M/F100LP, G235M/F170LP, and G395M/F290LP medium-resolution (R=500-1340) gratings[56], covering 0.7-5.2m. Three-shutter slitlets were employed, enabling a three-point nodding pattern to facilitate background subtraction. The total exposure time on sources 8595 and 9025 are the same at 3107s, while on source D36123 it is 18387s (6 higher).
The NIRSpec data processing followed the same methodology as used for other CEERS NIRSpec observations [27, 28, 57, 58, 59]. The NIRSpec data reduction was based on the STScI Calibration Pipeline [60] version 1.8.5 and the Calibration Reference Data System (CRDS) mapping 1041 for objects 8595 and 9025, and CRDS mapping 1029 for D36123. There are three main stages in the reduction process. Briefly, the calwebb_detector1 pipeline module was utilized to correct for the detector 1/f noise, subtract the bias and dark current, generate the count-rate maps (CRMs) from the uncalibrated images. There is an improvement in the jump step to correct the cosmic ray “snowball” [28]. The stage two of the pipeline reduction is to create two-dimensional (2D) spectra with a rectified trace and flat slope from the generated CRMs. The background subtraction and slit loss correction of objects D36123 and 8595 were considered at this stage, while the slit loss correction (pathloss) was not taken into account for object 9025 as it is poorly centered in the shutter. We verified that better overall of its spectral fit can be obtained if adopting the empirical aperture corrections directly, which is the last step of the absolute flux calibrations (as discussed later in a dedicated subsection) to complement any incorrect calibration during the reduction process. At the third stage, we made use of optimised apertures for the extraction of the one-dimensional (1D) spectra in all cases. For object D36123, we adopted a 4-pixel extraction aperture, considering the effects of the Micro-Shutter Assembly [61] (MSA) bar shadows on some upper pixels in 2D spectra. We also considered 3- and 5-pixel apertures, with negligible impact on the results. The flux error for the extracted spectra was calculated by the JWST pipeline using an instrumental noise. We rescaled all flux errors multiplying by a factor of 1.5 to account for correlations induced by the pipeline [27].
The spectrum of 9025, observed with 3 medium-resolution gratings, was calibrated through several steps. We first matched the 3 spectra scaling them by a constant term, determined using the overlapping regions, with variance weighting. This required a scaling of factors of 1.23 and 1.28 between G140M and G235M and between G235M and G395M, respectively. In the overlapping regions, the final coadded spectra were obtained by combining their weighted spectra and errors, ensuring that standard errors in the weighted mean over common wavelength ranges remained unchanged after resampling. For the purpose of increasing the continuum SNR to display the NIR features and speeding up the spectral fit computation, we finally resampled this medium-resolution spectrum to the one in the PRISM resolution. The average flux and the standard error of the mean in each PRISM wavelength bin were adopted.
4.2 Photometric measurements
[…]
| D36123 | 8595 | 9025 | |
| F150W | F160W | F150W | |
[…]
| n | 3.7130.025 | 3.020 0.823 | 2.0880.031 |
| b/a | 0.975 0.003 | 0.9000.090 | 0.7710.031 |
| PA | 13 | 4338 | -571 |
| Mag | 21.970.01 | 24.220.09 | 23.700.01 |
4.3 Aperture correction
Spectra reduced with the NIRSpec pipeline include an aperture correction term that is derived assuming a point-source scenario. To account for the actual shapes of our galaxies, we integrate the observed spectra through imaging filter bandpasses to obtain synthetic photometry (i.e., ), and compare to the actual photometry (i.e., ) measured from imaging. We are implicitly assuming here that no strong color gradient biases the in-slit photometry with respect to the integrated galaxy photometry. This point has been verified by building PSF-matched color images of our targets and verifying that no color-gradients were apparent. We thus derive the aperture correction as . The top panel of Fig.8 shows the aperture correction for object D36123 derived in this way. We adopt a linear function to fit the trend with wavelength, as shown by the red line. There is a weak wavelength dependence (slope 0.02). The green line presents the rescaled aperture correction derived from the forward-modeling tool for NIRSpec Multi-Object Spectroscopic data, namely MSAFIT [67], which accounts for the complex geometry of the target galaxy, point spread function and pixelation of the NIRSpec instrument. This MSAFIT line aligns with the red linear relation. Regardless of the relation (linear vs. MSAFIT) or when adopting the extremes in the 1 range in the linear relation (the grey shaded region of D36123 in Fig.8) to correct the spectrum, the best-fit parameters and spectral fits show minimal changes. Hence we decided to finally rescale the observed spectrum of D36123 by adopting the linear relation. The same process is applied to the other two objects 8595 and 9025. Their spectra are rescaled multiplying by a constant value of 2.28 and a linear relation with a slope of -0.07, respectively.
4.4 Spectral resolution
When comparing the observed spectra with spectral synthesis models we need to match their spectral resolution. For the observed spectra, we have to consider the instrumental broadening, as well as the intrinsic linewidths from the galaxies. For the template models, we need to estimate their intrinsic spectral resolution. We will discuss in turn these terms in the following.
The spectral resolution of the NIRSpec PRISM is relatively low. Based on the spectral sampling and for an unresolved source filling the slit, we can expect a resolution of 70 at 0.6m, declining to 30 at 1.1m, and rising again to 200 at 5m. For point/compact sources, the resolution of in-flight data could be 1.5-2.0 higher than that in the pre-launch approximation [67]. This would be helpful to identify some weak/narrower features, especially for galaxy D36123 where the SNR in the continuum is high. We adopt the MSAFIT [67] software to model the NIRSpec resolving power based on the actual galaxy spatial profile. The resolution modeled by MSAFIT software is consistent with that of a flat source in a slit width of 1.4 pixels (see green curve in the top-middle panel of Fig.9).
The intrinsic dispersion of our target early-type galaxies (ETGs) need to be accounted next. The velocity dispersion can be estimated using local stellar mass vs. velocity dispersion relations (i.e., -) relation [35]. We convert to the full width at half maximum (FWHM) and define (where FWHM). The intrinsic broadening of an ETG with a velocity dispersion 130km/s (a ballpark number for the typical masses in our study) corresponds to R1000 (see the grey line in the top-middle panel of Fig.9), quite negligible compared to the instrumental spread.
Finally, we estimated the template model’s resolution from their spectral sampling, adopting 2 spectral bin widths as their (Nyquist) sampling. The sampling (hence resolution) is changing in steps, higher in the optical than in the NIR. Redshifting will affect where these steps actually apply for the observed galaxies. The corresponding resolution in the observed frame for D36123 (assuming an average 1.077), is shown in the top-middle panel of Fig.9.
To match the resolution of the models with the observed spectrum, we need to account for all three impacting factors: prism dispersion (dominant term), the wavelength sampling of the spectral template models, and the velocity dispersion of the ETG. We add and subtract in quadrature the corresponding terms at fixed . The broadening needed to be applied to models to match the observed spectra is provided by the formula:
where refers to the model. The final smoothing functions applied to models are equivalent to resolutions among ranging from R 60 to 450 across observed wavelength range of 0.6–3.3m, as depicted in the top right panel of Fig.9. Beyond 3.3m we are limited by the models resolution and therefore no further smoothing is applied to them. We repeat similar procedures to obtain the appropriate smoothing kernels for objects 8595 and 9025, shown in the middle and bottom panels of Fig.9. For both, the prism resolution is based on 1.4-pixel elements, and the spectra in different models was resampled in every 2 spectral bin widths, similar to what is done for object D36123.
4.5 Spectral fit
To estimate the stellar mass, age, and SFH of a galaxy, we adopt a custom IDL routine [68] to fit the final corrected 1D spectrum of each galaxy. The best-fitting results are derived by comparing the 1D spectrum with a range of composite stellar population templates by -minimisation. These templates were generated by combining a grid of different SSP models assuming a delayed exponentially declining SFH (, i.e., delayed- model). The characteristic timescale and age variations are the same in different models, shown in Table 4. To identify the redshift, the Calzetti attenuation law [69] was adopted, and the stellar metallicity was left free with lower and upper limits dependent on the model. To reduce the computational cost, we run the spectral fit twice. First, we run the spectral fit in a large range of redshift with a low-resolution grid (). After the most probable redshift peak was identified, we narrowed down the redshift range with a high-resolution grid () in the second run. The redshift boundary in Table 4 is an example of a parametric grid for D36123 in the second run. For object D36123 with the highest SNR, the spectral fit was repeated varying the wavelength range (see Table 1): (1) with the full rest-frame wavelength (0.3-2m, see Fig 10); (2) with the optical only (m); (3) with the NIR only (m); and (4) with the full rest-frame wavelength range but excluding 0.5–1m (see Fig. 11). For the two fainter galaxies, the spectral fit was only done in the full rest-frame wavelength. Since the medium resolution grating spectrum of object 9025 does not cover too well the age-sensitive spectral region around rest-frame 4000 break, we additionally considered the photometry in HST F606W and F814W bands to better constrain the overall fit.
The best-fitting parameters with 1 errors[70] can be derived from the distribution of , which are listed in Table 2. Considering the large minimum reduced of object D36123, we derived the uncertainties (as listed in Table 1) scaling up all errors to obtain a 1.0. All stellar masses are reported for a Chabrier IMF [55], and are corrected for stellar losses and include remnants. Fig.10 shows the best-fit model for each galaxy using different models in red at the top of each panel, and the relative flux deviation at the bottom of each panel. The SNR-weighted averages of the relative deviations are labeled in the top-right corner of each bottom-panel. For object D36123 the models deviate from the spectrum at a few percent level (2.5% level for M13, and higher values for the other models). This is likely driven by systematics in the models (plus, possible residual calibration errors) which we can detect at the high SN ratio. Higher deviations correspond to observed features not matched by the models. For the other galaxies the relative flux deviations are larger and dominated by the noise in the data.
4.6 Additional details on models
The C09 models allow several user options and have been modified through the years. The models used in this work were generated using the latest FSPS version 3.2 by adopting a Chabrier IMF, the BaSTI isochrones and the MILES stellar library, without considering dust extinction. We correct for the expected stellar losses and include remnants. The C09 models include TP-AGB spectra, extrapolated blueward with simpler linear slopes (as advocated by Lancon & Wood 2000 [13]) and redward with the Aringer et al. (2009) [71] synthetic carbon star spectral library. The spectra of the oxygen-rich TP-AGB stars are extrapolated with the latest version of the PHOENIX stellar spectral library (the BT-SETTL library). For the TP-AGB isochrones, there is no shift adopted in the () vs. (). For crude understanding of the effect of TP-AGB stars in the NIR mass-to-light ratio we show 1 Gyr SSP models at solar metallicity from the different libraries, in Fig.12.
| Parameters [units] | Min | Max | Step |
| 1.06 | 1.1 | 0.001 | |
| Age [Gyr] | 0.1 | 3.5 | 0.1 |
| 0.001, 0.004, 0.009, 0.05 | |||
| [Gyr] | 0.1 | 1.0 | 0.1 |
| 1.0 | 3.0 | 1.0 | |
| Av [mag] | 0 | 2.1 | 0.1 (Calzetti) |
| 0.3 | 2.5 | 0.1 (BC03) | |
| 0.3 | 2.0 | 0.1 (C09) | |
| 0.3 | 2.2 | 0.1 (M13) | |
| 0.5 | 2.0 | 0.1 (M05) | |
The JWST NIRSpec data are available from the Mikulski Archive for Space Telescope (MAST; http://archive.stsci.edu), under program IDs 1345 and 2750. The CEERS JWST imaging data are available from MAST under program ID 1345. Reduced NIRCam data products from the CEERS team are available at https://ceers.github.io. The HST imaging data are available from the CANDELS survey at https://archive.stsci.edu/hlsp/candels, and the NIRCam/Ks imaging data are published in the 3D-HST survey at https://archive.stsci.edu/prepds/3d-hst/. We publicly release in a digital form of the calibrated spectrum of D36123 to enable tuning of future generations of spectral synthesis models (upon acceptance of this paper).
The JWST NIRSpec data were reduced using the JWST Pipeline (version 1.8.5, reference mapping 1041 and 1029; https://github.com.spacetelescope/jwst). The MSAFIT software is available at https://github.com/annadeg/jwst-msafit. The galfit software is published at https://users.obs.carnegiescience.edu/peng/work/galfit/galfit.html.
We thank Livia Origlia and Benne Holwerda for discussions. MD, PAH, SLF, JSK and CP acknowledge support from NASA through the STScI ERS award JWST-ERS-1345 and the JWST-GO-2750 award. CGG acknowledges support from CNES. CDE acknowledges funding from the MCIN/AEI (Spain) and the “NextGenerationEU”/PRTR (European Union) through the Juan de la Cierva-Formación program (FJC2021-047307-I). SL acknowledges the support by China Scholarship Council (CSC). This work is supported by the National Natural Science Foundation of China (No. 12192222, 12192220 and 12121003). This work is based on observations with the NASA/ESA/CSA James Webb Space Telescope obtained from the Mikulski Archive for Space Telescopes at the STScI, which is operated by the Association of Universities for Research in Astronomy (AURA), Incorporated, under NASA contract NAS5-03127.
SL and ED conceived this project, and selected and identified galaxies. SL led the analysis and together with ED, CM and MD wrote the manuscript. PAH and MD led the original observations and reduced the NIRSpec spectra. RG wrote the fitting code and with CDE helped with the spectral fitting procedures. All authors aided in the analysis and interpretation and contributed to the final manuscript.
The authors declare that they have no competing financial interests.